# Enumerate–Conjecture–Prove: Formally Solving Answer-Construction Problems in Math Competitions

### v2

@@ Abstract
Mathematical reasoning lies at the heart of artificial intelligence, underpinning applications in education, program verification, and research-level mathematical discovery. Mathematical competitions, in particular, presents two challenging problem types: theorem-proving, requiring rigorous proofs of stated conclusions, and answer-construction, involving hypothesizing and formally verifying mathematical objects. Large Language Models (LLMs) effectively generate creative candidate answers but struggle with formal verification, while symbolic provers ensure rigor but cannot efficiently handle creative conjecture generation. We introduce the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic method integrating LLM-based enumeration and pattern-driven conjecturing with formal theorem proving. We present ConstructiveBench, a dataset of 3,431 answer-construction problems in various math competitions with verified Lean formalizations. On ConstructiveBench dataset, ECP improves the accuracy of the answer construction from Chain-of-Thought (CoT) baseline of 14.54% to 45.06% with gpt-4.1-mini model. Moreover, combining with ECP’s constructed answers, state-of-the-art DeepSeek-Prover-V2-7B model generates correct proofs for 858 of the 3,431 constructive problems in Lean, achieving 25.01% accuracy, compared to 9.86% for symbolic-only baselines. Our code and dataset are publicly available at https://github.com/JackSun200312/ECP.
@@ 1 Introduction
Problems in math competitions broadly fall into two categories: theorem-proving and answer-construction tasks. In theorem-proving problems, a clear conclusion is stated, and the goal is a formal proof based on provided hypotheses. Conversely, answer-construction problems require generating specific mathematical objects (e.g., functions, numbers, sets) that satisfy given conditions, along with proofs of correctness. Representative examples of these problems appear in Table 1. Crucially, answer-construction problems pose the unique challenge of first identifying the correct candidate object or structure before verification.
@@
Table 1: Examples of competitional-level problems. The left shows an answer-construction problem requiring a candidate answer and proof. The right is a standard theorem-proving problem.
@@
Inspired by Mathematician George Pólya’s problem-solving methodology (Polya, 2014), which emphasizes systematically exploring simpler cases, identifying patterns, and rigorously proving general cases, we introduce the Enumerate–Conjecture–Prove (ECP) framework. Our framework combines the exploratory strengths of LLMs, which enumerate candidate answers through programmatic execution and generalize patterns to form conjectures, with the rigor of formal theorem proving methods that verify these conjectures. ECP thus effectively addresses the critical limitation of prior approaches by formally solving answer-construction problems, filling a significant gap in the mathematical reasoning landscape.
Our contributions are threefold: (i) a feedback-driven autoformalization pipeline for answer construction problems in math competition; (ii) ECP, a modular framework unifying LLM-driven code generation for enumeration, pattern-based conjecture for answer construction, and formal theorem proving; (iii) ConstructiveBench, a curated dataset of 3,431 competition-grade answer-construction problems with verified Lean formalizations and validated through experiments with LLMs and state-of-the-art theorem proving techniques.
@@ 2 Related Work
Formal methods such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008), CVC5 (Barbosa et al., 2022)) and first-order provers (e.g., Vampire (Kovács and Voronkov, 2013), E (Schulz, 2002)) are widely used for formal verification and logic reasoning. Interactive theorem provers (ITPs) like Isabelle (Paulson, 1994) and Lean (Moura and Ullrich, 2021) allow step-by-step formal proof development, often integrating automation via tools such as Sledgehammer (Böhme and Nipkow, 2010) and LeanCopilot (Song et al., 2024).
Recent neuro-symbolic systems integrate LLMs with formal proof environments. Goedel-Prover (Lin et al., 2025), DeepSeek-Prover series (Xin et al., 2024; Ren et al., 2025), and InternLM-Math (Ying et al., 2024b) leverage large-scale synthetic formal datasets and iterative training to generate Lean proofs at scale, achieving state-of-the-art results on benchmarks like miniF2F and PutnamBench. Domain-specific systems like AlphaGeometry (Trinh et al., 2024) and LIPS (Li et al., 2025) combine LLM-based intuition with symbolic reasoning to prove Olympiad geometry and inequality theorems, respectively.
@@
Recent work has explored several directions for enhancing the mathematical reasoning capabilities of LLMs. Prompting-based approaches (Wei et al., 2023; Zhang et al., 2025) use carefully crafted prompts to encourage intermediate steps or increase reasoning complexity. Fine-tuning strategies such as MAmmoTH (Yue et al., 2023) and MetaMath (Yu et al., 2024) improve performance by instruction-tuning on math datasets rich in intermediate rationales or diverse reformulations of mathematical problems. Reinforcement learning-based approaches  (Luo et al., 2025; Wang et al., 2023; Shao et al., 2024) improve reasoning by aligning step-by-step problem solving with reward signals. Tool-augmented frameworks like PAL (Gao et al., 2023) and Tora (Gou et al., 2023) offer an orthogonal enhancement by generating code as intermediate reasoning steps and outsourcing execution to Python interpreters or external computational tools. SATLM (Ye et al., 2023) extends this framework by using an Z3 symbolic solver to validate logic-based task specifications, though its primary focus is on formal logic tasks rather than mathematics.
@@
Benchmarks for evaluating LLM mathematical reasoning vary significantly in difficulty and format. Datasets like GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and MathQA (Amini et al., 2019) focus on grade-school and high-school level math with final-answer supervision. In contrast, OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024) include harder problems but still rely on string-matching for evaluation. Formal datasets such as miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), PutnamBench (Tsoukalas et al., 2024), and ProofNet (Azerbayev et al., 2023) allow machine-verifiable evaluation in Lean or HOL-Light. LeanWorkBook (Ying et al., 2024a) provides large-scale autoformalization from natural language, while Herald (Gao et al., 2024b) auto-informalize the Mathlib4 corpus back to natural language. PutnamBench remains actively maintained for recent versions of Lean.
@@ 3.1 Problem Formulation
Problems in math competitions can generally be classified into two categories: theorem-proving tasks and answer-construction tasks. In a theorem-proving task, the problem explicitly presents a set of hypotheses $P$ and a target conclusion $Q$, and the objective is to formally prove $P$ implies $Q$. Most prior benchmarks and research on AI for formal mathematics, such as MiniF2F (Zheng et al., 2022) and FIMO (Liu et al., 2023), focus on this task format, which aligns with traditional first-order theorem proving.
@@
The predicate $P:\alpha\rightarrow\text{Prop}$ that context variables satisfy.
@@
We demonstrate the definition by a simple straightforward example: “Given $a,b\in\mathbb{Z}$ with $b\neq 0,$ find all integers $x\in\mathbb{R}$ such that $ax=b$." Here, the context variables will be $(a,b)$ with type $\mathbb{Z}\times\mathbb{Z}$ and property $P(a,b):=b\neq 0$, the answer variable is $x$ with type $\mathbb{R}$, problem predicate is $Q(a,b,x):=ax=b$, and the answer function is $f(a,b)=\{\frac{b}{a}\}$.
In general, the task of answer-construction can be formulated as below in Lean’s type dependence theory:
$\exists f:\alpha\to\mathrm{Set}\ \beta,\ \forall(a:\alpha),\ P(a)\rightarrow\forall(b:\beta),\ Q(a,b)\leftrightarrow b\in f(a)$
@@
To ensure answers are not only correct but also semantically meaningful, we further require that constructed answers be expressed in canonical form, meaning they must not echo the problem statement’s syntax or structure, and must be in closed-form and algebraically simplified representation that generalizes across instances. Notably, Lean itself does not enforce canonical representations (e.g., $\mathbb{N}$, $\mathbb{N}$, or more complex yet equivalent expressions are all accepted), and certain LLM-based provers can trivially “cheat” by echoing the statement, e.g. setting $f(a,b)=\{x|ax=b\}$ in the above example or answering $\{f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}|x(f(x)+f(y))\geq(f(f(x))+y)f(y)\}$ for the example in Table 1. Such trivial substitutions lead to vacuously high proving accuracies if the answer-construction problem is handed straight to the prover. We therefore separate answer construction using tailored LLM prompts to avoid these trivial echoes before formal verification. Finally, the template and requirements for formalizing such answer-construction problems are incorporated into prompts for autoformalization described in Section 3.2 below.
@@ 3.2 Autoformalization Pipeline
Autoformalization translates informal, natural language mathematics into machine-verifiable formal statements in proof assistants such as Lean. Due to a scarcity of aligned informal-to-formal mathematical training data, LLMs still frequently fail at precise formalizations (Wu et al., 2022a), leading to syntactic and semantic errors.
To address this limitation, we implement a structured autoformalization pipeline that leverages LLMs, formal feedback from the Lean compiler, and targeted retrieval of formal definitions from external documentation sources, illustrated in Figure 1. The pipeline comprises three core components: (1) Autoformalizations with LLM (2) Lean knowledge retrieval and (3) error-guided feedback loops, and (3)
@@
We utilize specialized mathematical LLMs, including deepseek-math-7b-instruct (Shao et al., 2024) and advanced commercial models gpt-4o-mini, and deepseek-v3, for autoformalization. To better align these models with formalizations in high-school and undergraduate mathematics , we fine-tune deepseek-math-7b-instruct and gpt-4o-mini on a curated set of 791 paired informal–formal examples drawn from PutnamBench (Tsoukalas et al., 2024) and FIMO (Liu et al., 2023). Since deepseek-v3 does not yet support in-API fine-tuning, we apply a five-shot prompting strategy to guide its output.
@@
LLMs occasionally hallucinate non-existent definitions or misuse current Lean syntax due to their outdated training data or scarce training data in formalized statements. To mitigate this issue, we construct an external Lean knowledge base comprising 247,826 entries sourced from the official Lean 4.19.0 documentation and Mathlib (mathlib Community, 2020) repository via LeanDojo (Yang et al., 2023). After focusing on retaining only those definitions most relevant to high-school-level mathematical formalizations (Appendix A), the database is reduced to 44,477 entries. Each entry is represented by embeddings computed with Sentence Transformers (Reimers and Gurevych, 2019) to facilitate the retrieval of relevant definitions using FAISS (Douze et al., 2024). When an LLM encounters a missing or erroneous definition, the system retrieves the top candidate definitions using both FAISS similarity for embedding-based search and edit distance for fuzzy search, which enhances the formalization accuracy and reduces hallucination rates.
@@
Generated formalizations undergo immediate compilation checks in Lean, capturing syntactic errors, misuse of definitions, or semantic discrepancies. Each identified error triggers a structured feedback response, guiding the LLM toward correction in subsequent iterations. This iterative loop continues until formalizations compile successfully and satisfy semantic validation criteria evaluated by another LLM judge, which checks semantic equivalence with the original informal statement, proper usage of key definitions, and avoidance of trivial or uninformative answers. Additionally, we employ a sanity checker script to ensure the formalizations adhere to a unified format, which ensures consistency across the formalization process. Once this loop is complete, the subsequent enumerate–conjecture–prove stages take the formal statement as input and begin problem solving.
@@ 3.3 Enumerate–Conjecture–Prove Framework
ECP systematically integrates LLM’s capabilities of coding, mathematical reasoning with formal theorem proving within Lean proof assistant. Figure 2 visualizes the workflow of ECP applied to an example problem.
@@
Leveraging insights from the enumerated examples, a “Conjecturer" LLM attempts to solve the problem and construct a closed-form candidate answer in Lean through mathematical reasoning and pattern finding in enumerated answers. These conjectures propose candidate answers that are both syntactically simple and semantically meaningful, suitable for subsequent rigorous verification. We use a dedicated prompt template in this stage to enforce canonical answer formats and discourage trivial echoes of the original problem statement.
@@
In the final Prove stage, the constructed answer is substituted into the theorem statement, transforming the answer-construction problem into a standard theorem-proving task. The theorem is then verified using specialized LLM fine-tuned for theorem proving, including DeepSeek-Prover-V2 (Ren et al., 2025), Goedel-Prover (Lin et al., 2025), and Kimina-Prover (Wang et al., 2025). Additionally, we implement baseline symbolic methods as a sequence of automation tactics as follows:
@@
    simp: Simplifies expressions by rewriting
@@
    aesop (Limperg and From, 2023): Proves goal via best-first search proof engine.

### v3

@@ Abstract
Mathematical reasoning lies at the heart of artificial intelligence, underpinning applications in education, program verification, and research-level mathematical discovery. Mathematical competitions, in particular, present two challenging problem types: theorem-proving, which requires rigorous proofs of stated conclusions, and answer-construction, which involves hypothesizing and formally verifying mathematical objects. Large Language Models (LLMs) effectively generate creative candidate answers but struggle with formal verification, while symbolic provers ensure rigor but cannot efficiently handle creative conjecture generation. We introduce the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic method integrating LLM-based enumeration and pattern-driven conjecturing with formal theorem proving. We present ConstructiveBench, a dataset of 3,431 answer-construction problems in various math competitions with verified Lean formalizations. On the ConstructiveBench dataset, ECP improves the accuracy of answer construction from a Chain-of-Thought (CoT) baseline of 14.54% to 45.06% with the gpt-4.1-mini model. Moreover, combined with ECP’s constructed answers, the state-of-the-art DeepSeek-Prover-V2-7B model generates correct proofs for 858 of the 3,431 constructive problems in Lean, achieving 25.01% accuracy, compared to 9.86% for symbolic-only baselines. Our code and dataset are publicly available at https://github.com/JackSun200312/ECP.
@@ 1 Introduction
Mathematical competitions, in particular, present two challenging problem types: theorem-proving, requiring rigorous proofs of stated conclusions, and answer-construction, involving hypothesizing and formally verifying mathematical objects. In theorem-proving problems, a clear conclusion is stated, and the goal is a formal proof based on provided hypotheses. Conversely, answer-construction problems require generating specific mathematical objects (e.g., functions, numbers, sets) that satisfy given conditions, along with proofs of correctness. Representative examples of these problems appear in Table 1. Crucially, answer-construction problems pose the unique challenge of first identifying the correct candidate object or structure before verification.
@@
Table 1: Examples of competition-level problems. The left shows an answer-construction problem requiring a candidate answer and proof. The right is a standard theorem-proving problem.
@@
Inspired by mathematician George Pólya’s problem-solving methodology (Polya, 2014), which emphasizes systematically exploring simpler cases, identifying patterns, and rigorously proving general cases, we introduce the Enumerate–Conjecture–Prove (ECP) framework. Our framework combines the exploratory strengths of LLMs, which enumerate candidate answers through programmatic execution and generalize patterns to form conjectures, with the rigor of formal theorem-proving methods that verify these conjectures. ECP thus effectively addresses the critical limitation of prior approaches by formally solving answer-construction problems, filling a significant gap in the mathematical reasoning landscape.
Our contributions are threefold: (i) a feedback-driven autoformalization pipeline for answer-construction problems in math competitions; (ii) ECP, a modular framework unifying LLM-driven code generation for enumeration, pattern-based conjecture for answer construction, and formal theorem proving; (iii) ConstructiveBench, a curated dataset of 3,431 competition-grade answer-construction problems with verified Lean formalizations and validated through experiments with LLMs and state-of-the-art theorem-proving techniques.
@@ 2 Related Work
Formal methods such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008), CVC5 (Barbosa et al., 2022)) and first-order provers (e.g., Vampire (Kovács and Voronkov, 2013), E (Schulz, 2002)) are widely used for formal verification and logical reasoning. Interactive theorem provers (ITPs) like Isabelle (Paulson, 1994) and Lean (Moura and Ullrich, 2021) allow step-by-step formal proof development, often integrating automation via tools such as Sledgehammer (Böhme and Nipkow, 2010) and LeanCopilot (Song et al., 2024).
Recent neuro-symbolic systems integrate LLMs with formal proof environments. Goedel-Prover (Lin et al., 2025), the DeepSeek-Prover series (Xin et al., 2024; Ren et al., 2025), and InternLM-Math (Ying et al., 2024b) leverage large-scale synthetic formal datasets and iterative training to generate Lean proofs at scale, achieving state-of-the-art results on benchmarks like miniF2F and PutnamBench. Domain-specific systems like AlphaGeometry (Trinh et al., 2024) and LIPS (Li et al., 2025) combine LLM-based intuition with symbolic reasoning to prove Olympiad geometry and inequality theorems, respectively.
@@
Recent work has explored several directions for enhancing the mathematical reasoning capabilities of LLMs. Prompting-based approaches (Wei et al., 2023; Zhang et al., 2025) use carefully crafted prompts to encourage intermediate steps or increase reasoning complexity. Fine-tuning strategies such as MAmmoTH (Yue et al., 2023) and MetaMath (Yu et al., 2024) improve performance by instruction-tuning on math datasets rich in intermediate rationales or diverse reformulations of mathematical problems. Reinforcement learning-based approaches (Luo et al., 2025; Wang et al., 2023; Shao et al., 2024) improve reasoning by aligning step-by-step problem solving with reward signals. Tool-augmented frameworks like PAL (Gao et al., 2023) and Tora (Gou et al., 2023) offer an orthogonal enhancement by generating code as intermediate reasoning steps and outsourcing execution to Python interpreters or external computational tools. SATLM (Ye et al., 2023) extends this framework by using a Z3 symbolic solver to validate logic-based task specifications.
@@
Benchmarks for evaluating LLM mathematical reasoning vary significantly in difficulty and format. Datasets like GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and MathQA (Amini et al., 2019) focus on grade-school and high-school level math with final-answer supervision. In contrast, OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024) include harder problems but still rely on string-matching for evaluation. Formal datasets such as miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), PutnamBench (Tsoukalas et al., 2024), and ProofNet (Azerbayev et al., 2023) allow machine-verifiable evaluation in Lean or HOL-Light. LeanWorkBook (Ying et al., 2024a) provides large-scale autoformalization from natural language, while Herald (Gao et al., 2024b) auto-informalizes the Mathlib4 corpus back to natural language.
@@ 3.1 Problem Formulation
Problems in math competitions can generally be classified into two categories: theorem-proving tasks and answer-construction tasks. In a theorem-proving task, the problem explicitly presents a set of hypotheses $P$ and a target conclusion $Q$, and the objective is to formally prove that $P$ implies $Q$. Most prior benchmarks and research on AI for formal mathematics, such as MiniF2F (Zheng et al., 2022) and FIMO (Liu et al., 2023), focus on this task format, which aligns with traditional first-order theorem proving.
@@
The predicate $P:\alpha\rightarrow\text{Prop}$ that the context variables satisfy.
@@
We demonstrate the definition with a simple, straightforward example: “Given $a,b\in\mathbb{Z}$ with $b\neq 0$, find all integers $x\in\mathbb{R}$ such that $ax=b$." Here, the context variables are $(a,b)$ with type $\mathbb{Z}\times\mathbb{Z}$ and property $P(a,b):=b\neq 0$. The answer variable is $x$ with type $\mathbb{R}$, the problem predicate is $Q(a,b,x):=ax=b$, and the answer function is $f(a,b)=\{\frac{b}{a}\}$.
In general, the task of answer-construction can be formulated as follows in Lean’s dependent type theory:
$\exists f:\alpha\to\mathrm{Set}\ \beta,\ \forall(a:\alpha),\ P(a)\rightarrow\forall(b:\beta),\ Q(a,b)\leftrightarrow b\in f(a).$
@@
$\exists f:\alpha\rightarrow\beta,\ \forall(a:\alpha),\ P(a)\rightarrow\bigl{(}\forall(b:\beta),\,Q(a,b)\leftrightarrow b=f(a)\bigr{)}.$
When there are no context variables, the formulation reduces to $\exists f:\text{Set }\beta,\ \forall b:\beta,\ P(b)\leftrightarrow b\in f.$
To ensure answers are not only correct but also semantically meaningful, we further require that constructed answers be expressed in canonical form, meaning they must not echo the problem statement’s syntax or structure, and must be in a closed-form, algebraically simplified representation that generalizes across instances. Notably, Lean itself does not enforce canonical representations (e.g., $\mathbb{N}$, $\mathbb{N}$, or more complex yet equivalent expressions are all accepted), and certain LLM-based provers can trivially “cheat” by echoing the statement, e.g., setting $f(a,b)=\{x\mid ax=b\}$ in the above example or answering $\{f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}\mid x(f(x)+f(y))\geq(f(f(x))+y)f(y)\}$ for the example in Table 1. Such trivial substitutions lead to vacuously high proving accuracies if the answer-construction problem is handed straight to the prover. We therefore separate answer construction using tailored LLM prompts to avoid these trivial echoes before formal verification. Finally, the template and requirements for formalizing such answer-construction problems are incorporated into prompts for autoformalization described in Section 3.2 below.
@@ 3.2 Autoformalization Pipeline
Autoformalization translates informal, natural language mathematics into machine-verifiable formal statements in proof assistants such as Lean. Due to a scarcity of aligned informal-to-formal mathematical training data, LLMs still frequently fail to produce precise formalizations (Wu et al., 2022a), leading to syntactic and semantic errors.
To address this limitation, we implement a structured autoformalization pipeline that leverages LLMs, formal feedback from the Lean compiler, and targeted retrieval of formal definitions from external documentation sources, as illustrated in Figure 1. The pipeline comprises three core components: (1) autoformalizations with LLMs, (2) Lean knowledge retrieval, and (3) error-guided feedback loops.
@@
We utilize specialized mathematical LLMs, including deepseek-math-7b-instruct (Shao et al., 2024) and advanced commercial models gpt-4o-mini and deepseek-v3, for autoformalization. To better align these models with formalizations in high-school and undergraduate mathematics, we fine-tune deepseek-math-7b-instruct and gpt-4o-mini on a curated set of 791 paired informal–formal examples drawn from PutnamBench (Tsoukalas et al., 2024) and FIMO (Liu et al., 2023). Since deepseek-v3 does not yet support in-API fine-tuning, we apply a five-shot prompting strategy to guide its output.
@@
LLMs occasionally hallucinate non-existent definitions or misuse current Lean syntax due to their outdated or scarce training data in formalized statements. To mitigate this issue, we construct an external Lean knowledge base comprising 247,826 entries sourced from the official Lean 4.19.0 documentation and the Mathlib (mathlib Community, 2020) repository via LeanDojo (Yang et al., 2023). After focusing on retaining only those definitions most relevant to high-school-level mathematical formalizations (Appendix A), the database is reduced to 44,477 entries. Each entry is represented by embeddings computed with Sentence Transformers (Reimers and Gurevych, 2019) to facilitate the retrieval of relevant definitions using FAISS (Douze et al., 2024). When an LLM encounters a missing or erroneous definition, the system retrieves the top candidate definitions using both FAISS similarity for embedding-based search and edit distance for fuzzy search, which enhances formalization accuracy and reduces hallucination rates.
@@
Generated formalizations undergo immediate compilation checks in Lean, capturing syntactic errors, misuse of definitions, or semantic discrepancies. Each identified error triggers a structured feedback response, guiding the LLM toward correction in subsequent iterations. This iterative loop continues until formalizations compile successfully and satisfy semantic validation criteria evaluated by another LLM judge, which checks semantic equivalence with the original informal statement, proper usage of key definitions, and avoidance of trivial or uninformative answers. Additionally, we employ a sanity checker script to ensure the formalizations adhere to a unified format, ensuring consistency across the formalization process. Once this loop is complete, the subsequent enumerate–conjecture–prove stages take the formal statement as input and begin problem solving.
@@ 3.3 Enumerate–Conjecture–Prove Framework
ECP systematically integrates LLMs’ capabilities for coding and mathematical reasoning with formal theorem proving within the Lean proof assistant. Figure 2 visualizes the workflow of ECP applied to an example problem.
@@
Leveraging insights from the enumerated examples, a “Conjecturer" LLM attempts to solve the problem and construct a closed-form candidate answer in Lean through mathematical reasoning and pattern-finding in enumerated answers. These conjectures propose candidate answers that are both syntactically simple and semantically meaningful, suitable for subsequent rigorous verification. We use a dedicated prompt template in this stage to enforce canonical answer formats and discourage trivial echoes of the original problem statement.
@@
In the final Prove stage, the constructed answer is substituted into the theorem statement, transforming the answer-construction problem into a standard theorem-proving task. The theorem is then verified using specialized LLMs fine-tuned for theorem proving, including DeepSeek-Prover-V2 (Ren et al., 2025), Goedel-Prover (Lin et al., 2025), and Kimina-Prover (Wang et al., 2025). Additionally, we implement baseline symbolic methods as a sequence of automation tactics as follows:
@@
    simp: Simplifies expressions by rewriting.
@@
    aesop (Limperg and From, 2023): Proves the goal via a best-first search proof engine.